OpenStudy (anonymous):

4. In a 30°- 60° - 90° right triangle, the shorter leg has a length of 4. What is the length of the longer leg? 4 8 4/3 4/2

OpenStudy (anonymous):

help?

OpenStudy (raden):

shorter/longer = 1/2 4/x = 1/2 (assumed the longer leg is x solve for x

OpenStudy (anonymous):

oops i forgot to add the picture.

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

is that no. 2 ?

OpenStudy (anonymous):

4/2?

OpenStudy (raden):